# Shallow Flow Matching for Coarse-to-Fine Text-to-Speech Synthesis

### v1

@@ Abstract
We propose a shallow flow matching (SFM) mechanism to enhance flow matching (FM)-based text-to-speech (TTS) models within a coarse-to-fine generation paradigm. SFM constructs intermediate states along the FM paths using coarse output representations. During training, we introduce an orthogonal projection method to adaptively determine the temporal position of these states, and apply a principled construction strategy based on a single-segment piecewise flow. The SFM inference starts from the intermediate state rather than pure noise and focuses computation on the latter stages of the FM paths. We integrate SFM into multiple TTS models with a lightweight SFM head. Experiments show that SFM consistently improves the naturalness of synthesized speech in both objective and subjective evaluations, while significantly reducing inference when using adaptive-step ODE solvers. Demo and codes are available at https://ydqmkkx.github.io/SFMDemo/.
@@ 1 Introduction
FM has gained increasing attention due to its efficiency and high-quality generation in TTS. In conventional coarse-to-fine FM-based TTS models, coarse representations are used as conditions for the flow module. However, the generation still starts from pure noise, resulting in a suboptimal allocation of modeling capacity. Since the coarse representations already encode a significant portion of the overall semantic and acoustic structure, modeling the early stage from noise becomes redundant and contributes little to the quality of the final output. To deal with this issue, DiffSinger diffsinger has introduced a shallow diffusion mechanism on singing voice synthesis (SVS) and TTS. It uses a simple mel-spectrogram decoder, with the diffusion module starting generation at a shallow step based on the output of the simple decoder. Therefore, we extend the idea of shallow diffusion mechanism (with the name) on FM and propose shallow flow matching (SFM) for coarse-to-fine TTS models. During training, we construct intermediate states along the FM paths based on the coarse representations. To achieve this, we employ an orthogonal projection method to adaptively determine the corresponding time, define a principled construction approach, and formulate a single-segment piecewise flow. During inference, the generation starts from the intermediate state, skipping the early stages and focusing computation on the latter part of the flow. This leads to more stable generation and enhanced synthesis quality, and accelerates inference when using adaptive-step ordinary differential equation (ODE) solvers.
We validate the proposed SFM method on multiple coarse-to-fine TTS models, covering two mainstream FM architectures, U-Net unet and DiT dit , as well as two major types of semantic input, text sequences and speech tokens. Experimental results show that SFM consistently improves the naturalness of synthesized speech, as measured by both pseudo-MOS metrics and subjective evaluations. We also observe significant inference acceleration across various adaptive-step ODE solvers.
@@ 2.1 Flow matching
VF $\boldsymbol{u}_{t}$ induces a probability density path $p_{t}:[0,1]\times\mathbb{R}^{d}\rightarrow\mathbb{R}_{>0}$, which is a time-dependent probability density function (PDF). From time 00 to time $t$, the PDF of $\boldsymbol{x}$ is transported from $p_{0}(\boldsymbol{x}_{0})$ to $p_{t}(\boldsymbol{x}_{t})$ along $\boldsymbol{u}_{t}$. Chen et al. cnf proposed continuous normalizing flow (CNF) that models the $\boldsymbol{u}_{t}$ by a neural network $\boldsymbol{v_{\theta}}(\boldsymbol{x}_{t},t)$, where $\boldsymbol{\theta}$ are learnable parameters. A CNF reshapes a simple prior distribution $p_{0}$ into a complicated distribution $p_{1}$. Lipman et al. fm further proposed flow matching (FM), whose objective is $\mathcal{L}_{\mathrm{FM}}=\mathbb{E}_{t,p_{t}(\boldsymbol{x}_{t})}||\boldsymbol{v_{\theta}}(\boldsymbol{x}_{t},t)-\boldsymbol{u}_{t}(\boldsymbol{x}_{t})||^{2}$. Since appropriate $p_{t}$ and $\boldsymbol{u}_{t}$ are unknown, fm constructed probability paths conditioned on data sample $\boldsymbol{x}_{1}\sim q(\boldsymbol{x}_{1})$. Specifically, let $p_{0}(\boldsymbol{x}_{0})=\mathcal{N}(\boldsymbol{x}_{0}|\boldsymbol{0},\boldsymbol{I})$, $p_{1}(\boldsymbol{x}_{1})\approx q(\boldsymbol{x}_{1})$, the conditional probability path is $p_{t}(\boldsymbol{x}_{t}|\boldsymbol{x}_{1})=\mathcal{N}(\boldsymbol{x}_{t}|\boldsymbol{\mu}_{t}(\boldsymbol{x}_{1}),\sigma_{t}(\boldsymbol{x}_{1})^{2}\boldsymbol{I})$. The flow and VF are considered with the forms:
@@
Then, we can minimize the conditional flow matching (CFM) loss during training, which is proven by fm to be equivalent to minimizing FM loss $\mathcal{L}_{\mathrm{CFM}}=\mathbb{E}_{t,p_{t}(\boldsymbol{x}_{t})}\|\boldsymbol{v_{\theta}}(\boldsymbol{x}_{t},t)-\boldsymbol{u}_{t}(\boldsymbol{x}_{t}|\boldsymbol{x}_{1})\|^{2}$.
[…]
@@ 2.3 Flow matching-based TTS models
The FM module adopts an encoder-decoder structure. The Conformer-based conformer encoder encodes the speech tokens generated by the LLM into hidden states and linearly projects them to the same dimension as mel-spectrograms. These hidden states are then upsampled by a length regulator for token-spectrogram alignment and fed into the U-Net-based decoder with other conditions for FM training and inference. Therefore, the upsampled hidden states can be explicitly supervised to coarse mel-spectrograms.
@@ 3.1 Theorems
$\boldsymbol{x}_{\tau}$ =$=\begin{cases}\sqrt{(1-(1-\sigma_{\mathrm{min}})t_{m})^{2}-\sigma^{2}_{m}}\boldsymbol{x}_{0}+\boldsymbol{x}_{m},&\text{if }\Delta<1,\\ \frac{1}{\Delta}\boldsymbol{x}_{m},&\text{if }\Delta\geq 1,\end{cases}$ (7)
@@
$\boldsymbol{x}_{t}$ =$=\begin{cases}(1-\frac{t}{t_{m}})\boldsymbol{x}_{0}+\frac{t}{t_{m}}\boldsymbol{x}_{t_{m}},&\text{if }t<t_{m},\\ (1-\frac{t-t_{m}}{1-t_{m}})\boldsymbol{x}_{t_{m}}+\frac{t-t_{m}}{1-t_{m}}(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0}),&\text{if }t\geq t_{m},\end{cases}$ (9)
$\boldsymbol{u}_{t}$ =$=\begin{cases}\frac{1}{t_{m}}(\boldsymbol{x}_{t_{m}}-\boldsymbol{x}_{0}),&\text{if }t<t_{m},\\ \frac{1}{1-t_{m}}(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0}-\boldsymbol{x}_{t_{m}}),&\text{if }t\geq t_{m}.\end{cases}$ (10)
@@ 3.2 Proposed methods on coarse-to-fine FM-based TTS
$\mathcal{L}_{\mathrm{coarse}}=\mathbb{E}||\hat{\boldsymbol{X}}_{\boldsymbol{g}}-\boldsymbol{X}_{1}||^{2}.$ (11)
@@ 3.2.1 Orthogonal projection onto CondOT paths
$t_{\boldsymbol{h}}=\mathbb{E}[\frac{\mathrm{sg}[\hat{\boldsymbol{X}}_{\boldsymbol{h}}]\cdot\boldsymbol{X}_{1}}{\boldsymbol{X}_{1}\cdot\boldsymbol{X}_{1}}],\quad\sigma_{\boldsymbol{h}}^{2}=\mathbb{E}||\mathrm{sg}[\hat{\boldsymbol{X}}_{\boldsymbol{h}}]-t_{\boldsymbol{h}}\boldsymbol{X}_{1}||^{2},\quad\mathcal{L}_{\mu}=\mathbb{E}||\hat{\boldsymbol{X}}_{\boldsymbol{h}}-t_{\boldsymbol{h}}\boldsymbol{X}_{1}||^{2},$ (13)
@@ 3.2.2 Single-segment piecewise flow
where $\mathcal{S}$ is a time scheduler for the randomly sampled $t$.
@@
$\mathcal{L}_{\mathrm{CFM}}=\mathbb{E}_{t,p_{t}(\boldsymbol{X}_{t})}||\boldsymbol{v_{\theta}}(\boldsymbol{X}_{t},t)-\boldsymbol{U}_{t}||^{2},\quad\mathcal{L}_{\mathrm{SFM}}=\mathcal{L}_{\mathrm{coarse}}+\mathcal{L}_{t}+\mathcal{L}_{\sigma}+\mathcal{L}_{\mu}+\mathcal{L}_{\mathrm{CFM}}.$ (21)
@@ 3.2.3 Inference with SFM strength
[…]
@@ 4.2 Model implementations
SFM-c: In our SFM method, the coarse mel-spectrograms are not used as FM conditions. For ablation studies, we also evaluate a variant (denoted as SFM-c) that applies the SFM method and follows the conventional practice of using coarse mel-spectrograms as FM conditions.
@@ 4.3 Training and inference configurations
Matcha-TTS Vocos LJ Speech 800 128 4$\times$10-4 1 Euler
Matcha-TTS Vocos VCTK 800 128 2$\times$10-4 2 Euler
StableTTS Vocos VCTK 800 128 2$\times$10-4 10% 2 3 Dopri(5)
@@
CosyVoice-DiT Vocos LibriTTS 200 256 1$\times$10-4 10% 4 3 Dopri(5)
@@ 4.4.2 Subjective evaluation
CMOS: We use our proposed model with SFM as the reference system, and all other systems are compared against it. For each system, five utterances are selected and paired with corresponding utterances from the SFM model. Listeners are presented with utterance pairs and asked to rate the naturalness difference on a 7-point scale ranging from $-3$ to $+3$. A score of $+3$ indicates that the first utterance sounds much better than the second, while $-3$ indicates the opposite.
@@ 4.6 ODE solvers and speed analysis
For evaluations, we follow the official ODE solver settings of each model. For speed analysis with adaptive-step ODE solvers, we use the odeint from torchdiffeq torchdiffeq with various solvers. The default relative and absolute tolerances (1$\times$10-7 and 1$\times$10-9) are too strict, resulting in significantly longer inference time. Therefore, we adopt the settings used in StableTTS: 1$\times$10-5 for both tolerances. As shown in Table 2, we employ Huen(2), Fehlberg(2), Bosh(3), and Dopri(5) as adaptive-step solvers, where the number in () denotes the order.
We adopt real-time factor (RTF) as the metric for speed evaluation, which measures the ratio between the ODE solver inference time and the corresponding audio duration. To reduce the influence of runtime variance, we choose 100 utterances from each model’s validation set and run each solver five times, reporting the average RTF.
@@ 5.1 Evaluation results
SFM strength selection: As shown in Table 3 and the tables in Appendix D, using values of $\alpha>1.0$ significantly improves both pseudo-MOS scores and WER. The $\overline{\text{PMOS}}$, used as the selection criterion, tends to first increase and then decrease as $\alpha$ increases. This suggests that the adaptively determined $t_{\boldsymbol{h}}$ during training is typically small. As a result, the intermediate state constructed at inference with $\alpha=1.0$ corresponds to an early position on the CondOT paths. When Gaussian noise $\boldsymbol{X}_{0}$ is added at this stage, the resulting signal-to-noise ratio is low, making it difficult for the flow decoder to extract sufficient information and weakening the guidance during early sampling steps.
According to Eq. (13), $\hat{\boldsymbol{X}}_{\boldsymbol{h}}$ is supervised by the linearly down-scaled $t_{\boldsymbol{h}}\boldsymbol{X}_{1}$. This allows us to apply a larger $\alpha$ during inference to linearly scale up $\hat{\boldsymbol{X}}_{\boldsymbol{h}}$, thereby enhancing its guidance effect. However, due to estimation errors, increasing $\alpha$ also leads to a growing directional mismatch between $\alpha\hat{\boldsymbol{X}}_{\boldsymbol{h}}$ and $\boldsymbol{X}_{1}$. Meanwhile, the incorporated $\boldsymbol{X}_{0}$ decreases and results in a more deterministic sampling process. These factors ultimately reduce generation quality.
@@
Ground truth 4.380 3.964 4.241 3.566 1.000 $+0.22$
Reconstructed 4.085 3.739 4.208 3.472 0.993 $+0.12$
Baseline 4.186 3.692 4.282 3.308 0.971 $-0.48$
Ablated 4.217 3.763 4.311 3.355 0.972 $-0.27$
@@
Ground truth 3.999 3.562 3.986 1.534 1.000 $+0.16$
Reconstructed 3.819 3.246 3.977 1.666 0.985 $+0.08$
Baseline 4.008 2.978 3.870 1.534 0.939 $-0.31$*
Ablated 4.026 2.997 3.872 1.613 0.941 $-0.39$*
@@
Ground truth 3.999 3.562 3.986 1.534 1.000 $+0.48$*
Reconstructed 3.360 2.908 3.855 1.719 0.972 $+0.04$
Ablated 3.328 2.958 3.929 1.798 0.932 $-0.34$*
@@
Ground truth 4.136 3.262 4.345 3.180 1.000 $+0.19$ 3.40
Reconstructed 3.942 3.126 4.336 3.146 0.993 $-0.24$ 2.82*
Baseline 4.191 3.303 4.481 3.513 0.932 $-0.21$* 3.47
Ablated 4.183 3.369 4.487 3.578 0.932 $-0.14$ 3.58
@@
Ground truth 4.136 3.262 4.345 3.180 1.000 $+0.23$ 3.31
Reconstructed 3.322 2.855 4.211 3.144 0.989 $-0.12$ 2.86*
Ablated 3.499 3.086 4.316 3.614 0.936 $-0.31$* 3.15

### v2

@@ Abstract
We propose Shallow Flow Matching (SFM), a novel mechanism that enhances flow matching (FM)-based text-to-speech (TTS) models within a coarse-to-fine generation paradigm. Unlike conventional FM modules, which use the coarse representations from the weak generator as conditions, SFM constructs intermediate states along the FM paths from these representations. During training, we introduce an orthogonal projection method to adaptively determine the temporal position of these states, and apply a principled construction strategy based on a single-segment piecewise flow. The SFM inference starts from the intermediate state rather than pure noise, thereby focusing computation on the latter stages of the FM paths. We integrate SFM into multiple TTS models with a lightweight SFM head. Experiments demonstrate that SFM yields consistent gains in speech naturalness across both objective and subjective evaluations, and significantly accelerates inference when using adaptive-step ODE solvers. Demo and codes are available at https://ydqmkkx.github.io/SFMDemo/.
@@ 1 Introduction
FM has gained increasing attention due to its efficiency and high-quality generation in TTS. In conventional coarse-to-fine FM-based TTS models, coarse representations are used as conditions for the flow module. However, the generation still starts from pure noise, resulting in a suboptimal allocation of modeling capacity. Since the coarse representations already encode a significant portion of the overall semantic and acoustic structure, modeling the early stage from noise becomes redundant and contributes little to the quality of the final output. To deal with this issue, DiffSinger diffsinger has introduced a shallow diffusion mechanism on singing voice synthesis (SVS) and TTS. It uses a simple mel-spectrogram decoder, with the diffusion module starting generation at a shallow step based on the output of the decoder. Therefore, we extend the idea of shallow diffusion mechanism (with the name) on FM and propose shallow flow matching (SFM) for coarse-to-fine TTS models. During training, we construct intermediate states along the FM paths based on the coarse representations. To achieve this, we employ an orthogonal projection method to adaptively determine the corresponding time, define a principled construction approach, and formulate a single-segment piecewise flow. During inference, the generation starts from the intermediate state, skipping the early stages and focusing computation on the latter part of the flow. This leads to more stable generation and enhanced synthesis quality, and accelerates inference when using adaptive-step ordinary differential equation (ODE) solvers.
We validate the proposed SFM method on multiple coarse-to-fine TTS models, covering two mainstream FM architectures, U-Net unet and DiT dit . Experimental results show that SFM consistently improves the naturalness of synthesized speech, as measured by both pseudo-MOS metrics and subjective evaluations. We also observe significant inference acceleration across various adaptive-step ODE solvers.
@@ 2.1 Flow matching
[…]
@@
Then, we can minimize the conditional flow matching (CFM) loss during training, which is proven by fm to be equivalent to minimizing the FM loss $\mathcal{L}_{\mathrm{CFM}}=\mathbb{E}_{t,p_{t}(\boldsymbol{x}_{t})}\|\boldsymbol{v_{\theta}}(\boldsymbol{x}_{t},t)-\boldsymbol{u}_{t}(\boldsymbol{x}_{t}|\boldsymbol{x}_{1})\|^{2}$.
During inference, we use an ODE solver to solve the integral $\boldsymbol{x}_{\mathrm{pred}}=\boldsymbol{x}_{0}+\int_{0}^{1}\boldsymbol{v_{\theta}}(\boldsymbol{x}_{t},t)\,dt$.
@@ 2.3 Flow matching-based TTS models
The FM module adopts an encoder-decoder structure. The Conformer-based conformer encoder encodes the speech tokens generated by the LLM into hidden states and linearly projects them to the same dimension as mel-spectrograms. These hidden states are then upsampled by a length regulator for token-spectrogram alignment and fed into the U-Net-based decoder, along with other conditions, for FM training and inference. Therefore, the upsampled hidden states can be explicitly supervised to coarse mel-spectrograms.
@@ 3.1 Theorems
$\boldsymbol{x}_{\tau}$ $=\begin{cases}\sqrt{(1-(1-\sigma_{\mathrm{min}})t_{m})^{2}-\sigma^{2}_{m}}\boldsymbol{x}_{0}+\boldsymbol{x}_{m},&\text{if }\Delta<1,\\ \frac{1}{\Delta}\boldsymbol{x}_{m},&\text{if }\Delta\geq 1,\end{cases}$ (7)
@@
$\boldsymbol{x}_{t}$ $=\begin{cases}(1-\frac{t}{t_{m}})\boldsymbol{x}_{0}+\frac{t}{t_{m}}\boldsymbol{x}_{t_{m}},&\text{if }t<t_{m},\\ (1-\frac{t-t_{m}}{1-t_{m}})\boldsymbol{x}_{t_{m}}+\frac{t-t_{m}}{1-t_{m}}(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0}),&\text{if }t\geq t_{m},\end{cases}$ (9)
$\boldsymbol{u}_{t}$ $=\begin{cases}\frac{1}{t_{m}}(\boldsymbol{x}_{t_{m}}-\boldsymbol{x}_{0}),&\text{if }t<t_{m},\\ \frac{1}{1-t_{m}}(\boldsymbol{x}_{1}+\sigma_{\mathrm{min}}\boldsymbol{x}_{0}-\boldsymbol{x}_{t_{m}}),&\text{if }t\geq t_{m}.\end{cases}$ (10)
@@ 3.2 Proposed methods on coarse-to-fine FM-based TTS
[…]
@@ 3.2.1 Orthogonal projection onto CondOT paths
[…]
@@ 3.2.2 Single-segment piecewise flow
where $\mathcal{S}$ is an arbitrary time scheduler for the randomly sampled $t$.
@@
[…]
@@ 3.2.3 Inference with SFM strength
Substituting $\tilde{\boldsymbol{X}}_{\boldsymbol{h}}$, $\tilde{t}_{\boldsymbol{h}}$, and $\tilde{\sigma}^{2}_{\boldsymbol{h}}$ into Eq. (15) yields $\boldsymbol{X}_{\tilde{t}_{\boldsymbol{h}}}$. We can obtain the predicted results $\boldsymbol{X}_{\mathrm{pred}}$ by solving the integral $\boldsymbol{X}_{\mathrm{pred}}=\boldsymbol{X}_{\tilde{t}_{\boldsymbol{h}}}+\int_{\tilde{t}_{\boldsymbol{h}}}^{1}\boldsymbol{v_{\theta}}(\boldsymbol{X}_{t},t)\,dt$ with an ODE solver. For each model employing the SFM method, we determine the optimal $\alpha$ on the validation set. In most cases, the optimal $\alpha$ is relatively small, resulting in $\Delta=1$. The scaling factor $\frac{\alpha}{\Delta}$ has a theoretical upper bound of $\frac{1}{(1-\sigma_{\mathrm{min}})\hat{t}_{\boldsymbol{h}}+\hat{\sigma}_{\boldsymbol{h}}}$.
@@ 4.2 Model implementations
SFM-c: In our SFM method, the coarse mel-spectrograms are not used as FM conditions. For ablation studies, we also evaluate a variant (denoted as SFM-c) that not only applies the SFM method but also uses coarse mel-spectrograms as FM conditions.
@@ 4.3 Training and inference configurations
Table 1: Training and inference configurations of TTS models. $\beta$: CFG strength. $\dagger$}: 24,000 maximum frames.
@@
[…]
CosyVoice-DiT Vocos LibriTTS 200 256 $\times1$0-4 10% 4 3 Dopri(5)
@@ 4.4.2 Subjective evaluation
CMOS: We use our proposed model with SFM as the reference system, and all other systems are compared against it. For each system, five utterances are selected and paired with corresponding utterances from the SFM model. Listeners are presented with utterance pairs and asked to rate the naturalness difference on a 7-point scale ranging from 3-3 to +3+3. A score of +3+3 indicates that the first utterance sounds much better than the second, while 3-3 indicates the opposite.
@@ 4.6 ODE solvers and speed analysis
For evaluations, we follow the official ODE solver settings of each model. For speed analysis with adaptive-step ODE solvers, we use the odeint from torchdiffeq torchdiffeq with various solvers. The default relative and absolute tolerances ($\times1$0-7 and $\times1$0-9) are too strict, resulting in significantly longer inference time. Therefore, we adopt the settings used in StableTTS: $\times1$0-5 for both tolerances. As shown in Table 2, we employ Huen(2), Fehlberg(2), Bosh(3), and Dopri(5) as adaptive-step solvers, where the number in () denotes the order.
We adopt real-time factor (RTF) and number of function evaluations (NFE) as metrics for speed evaluation, where RTF measures the ratio between the ODE solver inference time and the corresponding audio duration, and NFE indicates how many times the solver queries the model. To reduce the influence of runtime variance, we choose 100 utterances from each model’s validation set and run each solver five times, reporting the average RTF. Note that the NFE is constant across runs.
@@ 5.1 Evaluation results
SFM strength selection: In Table 3 and the tables in Appendix D, $\tilde{t}_{\boldsymbol{g}}$ and $\tilde{\sigma}_{\boldsymbol{g}}$ are computed according to Eq. 22, and the reported values represent their means over all utterances in the validation sets. From these tables, it is evident that using values of $\alpha>1.0$ significantly improves both pseudo-MOS scores and WER. The $\overline{\text{PMOS}}$, which serves as the selection criterion, tends to increase initially and then decrease as $\alpha$ grows. This observation suggests that the adaptively determined $t_{\boldsymbol{h}}$ during training is generally small. As a result, the intermediate state constructed at inference with $\alpha=1.0$ corresponds to an early position on the CondOT paths. When Gaussian noise $\boldsymbol{X}_{0}$ is added at this stage, the resulting signal-to-noise ratio is low, making it difficult for the flow decoder to extract sufficient information and weakening the guidance during early sampling steps.
According to Eq. (13), $\hat{\boldsymbol{X}}_{\boldsymbol{h}}$ is supervised by the linearly down-scaled $t_{\boldsymbol{h}}\boldsymbol{X}_{1}$. This allows us to apply the $\alpha$ during inference to linearly scale up $\tilde{\boldsymbol{X}}_{\boldsymbol{h}}$ in Eq. 22, thereby enhancing its guidance effect. However, due to estimation errors, increasing $\alpha$ also leads to a growing distance between $\tilde{\boldsymbol{X}}_{\boldsymbol{h}}$ and $\tilde{t}_{\boldsymbol{h}}\boldsymbol{X}_{1}$. Meanwhile, the incorporated $\boldsymbol{X}_{0}$ decreases and results in a more deterministic sampling process. These factors ultimately reduce generation quality.
@@
Ground truth 4.380 3.964 4.241 3.566 1.000 +0.22+0.22
Reconstructed 4.085 3.739 4.208 3.472 0.993 +0.12+0.12
Baseline 4.186 3.692 4.282 3.308 0.971 0.48-0.48
Ablated 4.217 3.763 4.311 3.355 0.972 0.27-0.27
@@
Ground truth 3.999 3.562 3.986 1.534 1.000 +0.16+0.16
Reconstructed 3.819 3.246 3.977 1.666 0.985 +0.08+0.08
Baseline 4.008 2.978 3.870 1.534 0.939 0.31-0.31*
Ablated 4.026 2.997 3.872 1.613 0.941 0.39-0.39*
@@
Ground truth 3.999 3.562 3.986 1.534 1.000 +0.48+0.48*
Reconstructed 3.360 2.908 3.855 1.719 0.972 +0.04+0.04
Ablated 3.328 2.958 3.929 1.798 0.932 0.34-0.34*
@@
Ground truth 4.136 3.262 4.345 3.180 1.000 +0.19+0.19 3.40
Reconstructed 3.942 3.126 4.336 3.146 0.993 0.24-0.24 2.82*
Baseline 4.191 3.303 4.481 3.513 0.932 0.21-0.21* 3.47
Ablated 4.183 3.369 4.487 3.578 0.932 0.14-0.14 3.58
@@
Ground truth 4.136 3.262 4.345 3.180 1.000 +0.23+0.23 3.31
Reconstructed 3.322 2.855 4.211 3.144 0.989 0.12-0.12 2.86*
Ablated 3.499 3.086 4.316 3.614 0.936 0.31-0.31* 3.15